OpenStudy (anonymous):

use substitution so solve 3x+4y=-3 x+2y=-1

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

I can help you out on this one buddy. :)

OpenStudy (anonymous):

please do

OpenStudy (anonymous):

Firstly, you isolate y in your first equation in the goal of replacing it in your second equation.

OpenStudy (anonymous):

4y=-3x-3 y=(-3x-3)/4

OpenStudy (anonymous):

Then you replace the new y value into the second equation.

OpenStudy (anonymous):

Try that and mention me when you think you found an answer.

OpenStudy (dtan5457):

I got you. Line them up together and use the elimination method. 3x+4y=-3 x+2y=-1 Multiply the last one by 3. 3x+4y=-3 3x+6y=-3 Subtract. -2y=0 Therefore, y=0. Then, to get x, put the y which is 0, to the 2nd equation. x+2(0)=-1 x=-1

OpenStudy (anonymous):

thanks guys